To what temperature must a balloon, initially at 47 C and 3.00 L, be heated in order to have volume of 7.00 L ?
1
Expert's answer
2018-10-24T08:52:09-0400

To what temperature must a balloon, initially at 47 C and 3.00 L, be heated in order to have volume of 7.00 L ?

Answer:

PV=nRT

P=const

R=const

n=?

n=V/T = 3 / 320 = 0,009375

T2 = V2 / n = 7 / 0.009375 = 746,67 K or 473.67 C